OpenStudy (scarlettfarra2000):

Find the coordinates of the vertex and the intercepts of the following quadraction function. When necessary, approximate the x – intercepts to the nearest tenth. r(x)=-3x^2-3x-2

OpenStudy (scarlettfarra2000):

A quadratic function is a function of the form f(x)= ax^2+bx+c where a, b, and c are real numbers anda≠0. For the given functions a=-3 b=-3 and c=-2 The vertex occurs at x=-b/2a to find the x- coordinate of the vertex of the given functions, substitute the values of a and b into the expression and simplify Let a=-3 and b=-3 simplify x=-b/2a=-3/2(-3) =-3/6 The vertex has an x – coordinate of -3/6. To find the y – coordinate, we evaluate r(-3/6) \[r(-\frac{ 3 }{ 6})=-3(-\frac{ 3 }{ 6 })-3(-\frac{ 3 }{ 6 })-2=\frac{ 11 }{ 36}\] The vertex would be \[\frac{ 11 }{ 36},-\frac{ 3 }{ 6 }\]

OpenStudy (scarlettfarra2000):

@Atrineas

OpenStudy (atrineas):

Idk how to use the equation thing xD but yeah i think that would be it but flip your vertex around cause (x,y) and -3/6 is your x when 11/36 is y

OpenStudy (scarlettfarra2000):

Nope

OpenStudy (atrineas):

@TheSmartOne please help

OpenStudy (scarlettfarra2000):

@jim_thompson5910 if your not busy could you please help me

OpenStudy (scarlettfarra2000):

I can't figure out where I did something wrong

OpenStudy (atrineas):

@satellite73 can you help her if you have the time please vertex needs to be in fraction and simplified if possible

OpenStudy (fortytherapper):

I have not did the math, but.... For the first part of plugging in -3/6 \[-3(\frac{ -3 }{ 6})-3(\frac{- 3 }{ 6 })-2\] Did you square the first part?

OpenStudy (scarlettfarra2000):

to answer you question @FortyTheRapper I did but don't know if I did it right or not

OpenStudy (fortytherapper):

I got something different: -3/6 is the same thing as -1/2, which is the same thing as .5 \[(x,y) \rightarrow (-\frac{ 1 }{ 2 },y)\] So I would evaluate \[y = -3(0.5)^2-3(0.5)-2\]

OpenStudy (scarlettfarra2000):

Here's the whole problem because I'm thinking we're suppose to get fractions that's why I've got it wrong the last three or four times (And ignore what I put in the box it's wrong anyways)

OpenStudy (fortytherapper):

It says integers or fractions, so either will work Also, -3/6 should be first, since that's your x value

OpenStudy (scarlettfarra2000):

Okay

OpenStudy (scarlettfarra2000):

But would 0.5 be the y?

OpenStudy (fortytherapper):

It would be the x, and negative. As you said in the first response: "The vertex occurs at x=-b/2a to find the x- coordinate of the vertex of the given functions, substitute the values of a and b into the expression and simplify"

OpenStudy (scarlettfarra2000):

right

OpenStudy (scarlettfarra2000):

Now how do we find the y?

OpenStudy (fortytherapper):

We plug .5 into the equation that was given

OpenStudy (scarlettfarra2000):

\[-3(0.5)-3(0.5)-2 = 0.15(2)-0.15(2)\] because you still have to add 2 both times right?

OpenStudy (fortytherapper):

Remember for the first -3(0.5), you have to square the 0.5 first then multiply by -3

OpenStudy (scarlettfarra2000):

So it'd become \[-3(0.25)-3(0.25)-2=-0.75-0.75-2?\]

OpenStudy (fortytherapper):

Almost got it But you squared the second 0.5, when you should have just squared the first 0.5, which means you have this so far: \[y=-0.75-3(0.5)-2\] After you calculate the middle term, we can just do simple subtraction and get a Y value

OpenStudy (scarlettfarra2000):

okay so it'd be \[(-\frac{ 3 }{ 6 }, 1.5)\]

OpenStudy (fortytherapper):

No, because \[-0.75-3(0.5) -2 \neq1.5\]

OpenStudy (fortytherapper):

First, do 3(0.5)

OpenStudy (fortytherapper):

Then you can subtract after finding that value

OpenStudy (scarlettfarra2000):

the answers been in front of my face the whole time haven't it?

OpenStudy (fortytherapper):

Somewhat, we just have to multiply 3 and 0.5 in order to get to the last step

OpenStudy (scarlettfarra2000):

1.5

OpenStudy (fortytherapper):

Correct, now we can find the answer \[y = -0.75-1.5-2\]

OpenStudy (scarlettfarra2000):

-2.75 or -0.5?

OpenStudy (fortytherapper):

Subtract it one step at a time I would do -1.5-2 first, then subtract 0.75 from that answer to get the y value

OpenStudy (scarlettfarra2000):

you get 4.25

OpenStudy (fortytherapper):

Yep good job, and that's negative. So the answer is (-0.5, -4.25)

OpenStudy (scarlettfarra2000):

Okay let me see and there's going to be two more parts after this one

OpenStudy (scarlettfarra2000):

It's says it's wrong too

OpenStudy (fortytherapper):

I see how it is, I forgot the negative. Mb, lemme fix it real quick for you. Sorry

OpenStudy (scarlettfarra2000):

it's okay

OpenStudy (fortytherapper):

Since you understand it all, I'll just say \[y = -3(-0.5)^2-3(-0.5) - 2 = -0.75 + 1.5-2 = 0.75\] So the answer should be \[(-0.5,0.75)\]

OpenStudy (scarlettfarra2000):

No it's wrong again the answer is \[(-\frac{ 1 }{ 2 }, -\frac{ 5 }{ 4 })\]

OpenStudy (fortytherapper):

Ima have to disagree with MathXL on this one, because I don't see how it can be that

OpenStudy (scarlettfarra2000):

I know how you feel well it's okay I got the rest right now I've got to restart a whole new problem because I can't go on until all problems are 100%

OpenStudy (fortytherapper):

I'll rework this one and see what the problem is, if there even is one

OpenStudy (scarlettfarra2000):

you don't have to it's fine

OpenStudy (fortytherapper):

I must, I'm one of those people T_T

OpenStudy (scarlettfarra2000):

Alright it's mostly up to you but thank you for all your help